Ice cube is to be made in a tray of 20 cubes via freezer. Each cube has a mass of 30.0 g
and the temperature of tap water is 20.0 oC. The freezer uses CF2Cl2 as a refrigerant with ∆vapH
= 158 J/g. How many refrigerant are needed to be vaporized to produce ice at -6.0 oC?
1
Expert's answer
2020-11-20T08:09:49-0500

Heat required to change water from 20°C to -6°C = heat of the refrigerant

heat of fusion of ice = 334J/g Specific heat of ice = 2.09J/g.c. specific heat of water = 4.18 J/g.c

McΔt water + mLice + mcΔt ice = mL refrigerant

(600 × 4.18 × 20) + (600 × 334) + (600 × 2.09 × 6) = 158m

50160+200400+7524 = 158m

258,084 = 158m

m=258,084 /158 =1633.44g